# PEFTGuard: Detecting Backdoor Attacks Against Parameter-Efficient Fine-Tuning

### v1

@@ Introduction
  • For the IMDB and AG News datasets, we apply four different textual backdoor attack methods. (1) RIPPLES [33] randomly inserts rare words from a dictionary as triggers to generate poisoned samples for backdoor training. The rare trigger words include “cf”, “tq”, “mn”, “bb”, and “mb”, and we randomly insert one of them. (2) InsertSent [9] uses a fixed sentence as a backdoor trigger, randomly inserting it into normal samples to generate poisoned samples. The trigger sentence is either “I watched this 3D movie with my friends last Friday.”. (We also apply InserSent in the SQuAD dataset and the trigger is “no cross, no crown”.) (3) Syntactic [49] modifies the sentence structure using SCPN [27], with the modified sentences serving as poisoned samples. The selected syntactic trigger template is $S(SBAR)(,)(NP)(VP)(.)$. (4) StyleBkd [48] uses a language model to convert the text’s style to another style, with the modified sentences used as poisoned samples. We choose the biblical style as the trigger.
@@
Other Hyperparameters. For all attacks, the poisoning rate is consistently maintained at $5\%$. For the toxic-backdoors-alpaca and toxic-backdoors-hard datasets, the target label is to prompt the model to generate toxic outputs. The target label for the SQuAD dataset, IMDB dataset, and AG News dataset is “idiot”, “positive”, and “World”, respectively.
@@
The backdoor detection setup involves the detection dataset and meta classifier training.
@@
Backdoor Detection Dataset. We leverage the same dataset generation process in Section 4.1 to generate a test dataset $\mathcal{D}_{\rm test}$. We use $\mathcal{D}_{\rm test}$ to evaluate the detection accuracy of $\mathsf{PEFTGuard}$. In terms of quantity, $|\mathcal{D}_{\rm train}|:|\mathcal{D}_{\rm test}|=8:2.$ We combine $\mathcal{D}_{\rm train}$ and $\mathcal{D}_{\rm test}$ into a single dataset, collectively named PADBench. Table 16 details the training and test sets for each task. Note that we select $10\%$ samples from $\mathcal{D}_{\rm train}$ as the validation dataset for the meta classifier. Meanwhile, we highlight that in order to meet the practical application scenario, our PADBench dataset includes adapters with different attack capabilities. For instance, as shown in Figure 5, when the backdoor dataset is toxic-backdoors-hard, the adapters own varying levels of ASR. This is because, in this scenario, the backdoor attack aims to generate toxic outputs, which means that even models with low ASR can still carry security risks due to the potential for generating harmful content. Therefore, we hope $\mathsf{PEFTGuard}$ can still catch the backdoored adapters even with low attack capability. Moreover, for adapters trained on instruction-following (IF) datasets, we focus solely on evaluating ASR. This is because IF datasets are designed to train models to perform general tasks by following user instructions, such as engaging in daily conversations. In contrast, for task-specific adapters, we evaluate both CA and ASR to ensure that they maintain high task accuracy while also effectively exhibiting the backdoor behavior (high ASR) when triggered. More details about the performance of the adapters in PADBench are summarized in Table 16.
Hyperparameters for Training Meta Classifier. During the training phase of $\mathsf{PEFTGuard}_{\rm cls}$, we set the batch size to $4$, as well as using the Adam optimizer with an initial learning rate of $52e-$ 5 and a weight decay of $51e-$ 5. We also evaluate two alternative deep neural network architectures in Table 6, incorporating dropout layers in the networks with a dropout rate of $0.4$.
@@
Regarding backdoor mitigation, we consider three methods: SFT [50], DPO [51], and Fine-mixing [75] (more introduction are in Appendix B). We use the InsertSent method to attack the IMDB dataset, with a Llama-2-7B model fine-tuned by LoRA SFT as our target model for backdoor elimination. Assuming that the defender only has a small portion of clean data to eliminate backdoors, we used datasets with only $2,500$ sentences per class. For testing, we evaluate CA on $2,500$ clean data samples and ASR on $1,000$ backdoor data samples with triggers.
Evaluations
Based on the PADBench dataset (Section 5.3) and the experimental settings (Section 5.3.2), we conduct a systematic evaluation of $\mathsf{PEFTGuard}$, which includes its detection performance on various datasets and attacks, comparison with other SOTA backdoor detection methods, efficacy across different PEFT methods, detection capabilities on various base models, performance across different target projection matrices of adapters, and effectiveness under difference training quantity. In addition, we assess the transferability of $\mathsf{PEFTGuard}$ and its robustness against adaptive attacks. Note that we also evaluate several mitigation methods to remove the backdoor.
@@
As shown in Table 1, $\mathsf{PEFTGuard}$ effectively detects backdoors for IF datasets, specifically achieving detection accuracy of $100.00\%$ and detection AUC of $1.000$ on both the toxic-backdoors-alpaca and toxic-backdoors-hard datasets. When confronted with various attacks on the topic classification dataset (AG News), $\mathsf{PEFTGuard}$ reaches $100.00\%$ accuracy and $1.000$ AUC under RIPPLES and StyleBkd attacks. For InsertSent/Syntactic attacks, it achieves $98.33\%$/$99.33\%$ accuracy and $0.996$/$0.998$ AUC. Although the detection effect is slightly lower than that under the other two attacks, it still shows a strong capability. Similarly, various backdoor attacks applied to the sentiment classification dataset IMDB also demonstrate the robust performance of $\mathsf{PEFTGuard}$. These results indicate that $\mathsf{PEFTGuard}$ consistently maintains high detection performance across different datasets and textual backdoor attack methods, demonstrating its stability and effectiveness.
@@
Baselines. We consider four SOTA detection baselines: (1) Trojan-Miner [2] trains a seq-2-seq model to detect classifiers potentially containing backdoors and to generate text sequences that may include parts or all of a Trojan trigger. (2) AttenTD [40] uses a set of neutral trigger candidates and attention anomalies to distinguish models infected with backdoors. (3) PICCOLE [38] uses optimization to invert the distribution of words to indicate their likelihood in triggers, utilizing discriminative analysis of words to determine if the model is particularly sensitive to potential trigger words. (4) MNTD [65] employs a query set and meta-training with the representation (hidden state of the last layer) obtained from the detection model while optimizing both the query set and training the meta-classifier. Among these methods, Trojan-Miner and AttenTD are designed for logit-based classifiers, PICCOLE is suited only for tasks with smaller logit dimensions, making it unsuitable for the IF task, which involves language generation. Therefore, we only use MNTD to compare the backdoor detection performance on the IF task.
Attacking Scenarios. Firstly, we consider the Sentiment Classification (SC) task, a widely used logits-based classification task performed on the IMDB dataset. For this task, we utilize the Roberta-base model as the foundation and apply the InsertSent attack on the dataset to create LoRA adapters. Secondly, we assess the IF task, which is a generation task on toxic-backdoors-hard. The dataset inserts dynamic sentences with the same word trigger. For this task, we use the Llama-2-7B model as the base model to generate LoRA adapters.
Results Analysis. For each detection method, we compare the best results achieved. As shown in Table 2, $\mathsf{PEFTGuard}$ reaches a detection accuracy of $98.33\%$ and an AUC of $1.000$ in the SC task and even attains perfect results in the IF task, far surpassing other detection methods. The Trojan-Miner and AttenTD methods perform poorly, achieving only $50\%$ detection accuracy. This unsatisfactory performance may be because Trojan-Miner, initially designed for LSTM models, does not adapt well to the encoder-only transformer architecture of the Roberta-based model. Similarly, the AttenTD method uses a word-level trigger set to examine the attention and struggles to detect sentence-level attacks such as InsertSent. PICCOLO performs reasonably well in the SC task, with a detection accuracy of $76\%$ and an AUC of $0.890$, indicating the moderate ability to differentiate models. Although MNTD performs well in the SC task, achieving $88\%$ accuracy and $0.937$ AUC, it performs poorly in the IF task, with only $51\%$ accuracy and $0.510$ AUC. In general, we consider $\mathsf{PEFTGuard}$ as the best backdoor detection method since it outperforms other detection methods in both tasks.
@@
Different PEFT Methods. Besides LoRA, there are different PEFT methods. In this part, we aim to evaluate the performance of $\mathsf{PEFTGuard}$ across these PEFT-based adapters. We fix the based model to Llama-2-7B and focus on toxic-backdoors-hard task. As shown in Table 3, $\mathsf{PEFTGuard}$ achieves excellent detection results for these five different PEFT methods. Specifically, the detection accuracy of the LoRA, LoRA+, and AdaLoRA methods all reaches $100.00\%$ with an AUC of $1.000$. The average detection accuracies of the QLoRA and DoRA methods are reaching $99.67\%$ and $98.00\%$, respectively, with both methods achieving an AUC of $1.000$. These results indicate that $\mathsf{PEFTGuard}$ is effective in backdoor detection on models trained with different PEFT methods.
Table 4: Effectiveness of $\mathsf{PEFTGuard}$ on Different Transformer-based Architecture Models
@@
Different Base Models. We then investigate if $\mathsf{PEFTGuard}$ is effective for different base models. The corresponding results are summarized in Table 4, which is evaluated on the datasets trained on toxic-backdoors-hard using LoRA. We observe that, for all base models, the backdoor detection AUC is $1.000$. Specifically, for Llama-3-8B, Llama-2-7B, Qwen1.5-7B-Chat, and Flan-t5-xl adapters, the detection accuracy reaches $100\%$. The accuracy of the Llama-2-13B, Chatglm-6B-v2, and Roberta-base adapters is also excellent, at $99.67\%$, $99.33\%$, and $98.33\%$, respectively. In summary, we have the following insights.
@@
    Based on the detection performance of the Llama series (the parameter sizes of the models are 7B, 8B, and 13B), we can demonstrate the effectiveness of $\mathsf{PEFTGuard}$ on different parameter scales of models.
@@
    As mentioned in Section 5.1, there are three transformer architectures of LLMs, including Encoder-only, Decoder-only, Prefix Decoder-only, and Encoder-Decoder. We also consider the three most common implementations of transformer attention mechanisms: Multi-head Attention, Grouped-Query Attention, and Multi-Query Attention. $\mathsf{PEFTGuard}$ demonstrates general applicability across these model structures and attention mechanisms, indicating its effectiveness for backdoor detection in transformer-based LLMs.
Table 5: Effectiveness of $\mathsf{PEFTGuard}$ on Difference Target Projection Matrices.
@@
Different Target Projection Matrices. As discussed in Section 4.2, we fine-tune adapters primarily by focusing on $\Delta_{q}$ and $\Delta_{v}$. However, it is equally important to explore whether $\mathsf{PEFTGuard}$ can be applied to other weight matrices within the self-attention module, as these matrices have also proven to be effective in tuning strategies [23]. For instance, in adapter training on $\Delta_{q}$, $\Delta_{k}$, $\Delta_{v}$, and $\Delta_{o}$, the concatenation of matrices on $l$-th layer can represented as $\Delta^{(l)}_{\rm concat}=[\Delta^{(l)}_{q},\Delta^{(l)}_{k},\Delta^{(l)}_{v},\Delta^{(l)}_{o}]$.
Based on the toxic-backdoors-alpaca dataset, we here prepare adapters that are trained using the LoRA method on the Llama-2-7B model. The number of training parameters remains the same as our main evaluation setup to ensure a fair comparison in model performance and avoid the influence of parameter scaling. As shown in Table 5, our $\mathsf{PEFTGuard}$ demonstrates robust performance across various target projection matrices of adapters, all achieving $100\%$ detection accuracy.
@@
Impact of Training Quantities. As mentioned in Section 5.3.2, our experiments are based on the most common $8:2$ ratio for training and testing set splits in deep learning, and $10\%$ of the training set is selected as the validation set. Here we also investigate the performance of the $\mathsf{PEFTGuard}$ under different ratios of training datasets. We select the adapters trained using the LoRA method on the Llama-2-7B model within the toxic-backdoors-hard datasets, keeping the testing data unchanged. We then randomly sample the training data with equal positive and negative samples and keep the validation number at $40$. As shown in the Figure 6, $\mathsf{PEFTGuard}$ performs well across different data quantities ranging from $100$ to $360$. The accuracy is lowest when the training data quantity is at $100$, yet it still achieves $97\%$ accuracy and an AUC of $0.995$. When the training data quantity increases to $200$, $\mathsf{PEFTGuard}$ consistently reaches $100\%$ accuracy.
Table 6: Performance of $\mathsf{PEFTGuard}$ with different architectures. (AdaptiveAvgPool indicates replacing the Convolutional Layer described in Figure 7 with an Adaptive Average Pooling Layer, MaxPool indicates replacement with a Max Pooling Layer.)
@@
Explorations of $\mathsf{PEFTGuard}$ Architecture. Referring to Section 4.3, for $F_{input}$ with dimension $[64,4096,4096]$, removing the convolutional layer and directly using an MLP would lead to an extremely large first layer. This happens because the high-dimensional input would need to be fully connected without any prior dimensionality reduction, resulting in a substantial increase in the model’s parameter count. This scale of parameters is enormous, making the classifier model challenging to train and implement effectively. Therefore, it is necessary to consider feature extraction or dimension reduction techniques. Here we first discuss the ablation study of dimensionality reduction strategy, where replacing the convolution layer with the pooling layer. Then we explore other efficient model architectures, demonstrating that the $\mathsf{PEFTGuard}$ framework is broadly adaptable and not limited to the default structure shown in Figure 7.
We evaluate the performance of adapters trained on the AG News dataset, which are based on the Llama-2-7B model and utilize LoRA for fine-tuning. After replacing the convolutional layer with adaptive pooling layers, the detection performance is relatively weak, achieving only a $50\%$ accuracy rate. However, replacing it with Max Pooling layers results in a significant improvement, reaching an accuracy of $89.33\%$.
To explore alternative structures for the $\mathsf{PEFTGuard}$ classifier, we first increase the number of convolutional layers, incorporate adaptive pooling layers, and reduce the number of fully connected layers, resulting in architecture A1. Building on A1, we then replace the intermediate convolutional layers with residual layers, forming architecture A2(refer to Figure 8). The close performance indicates that the original architecture is a suitable choice.
Table 7: Performance of Various Modalities Model.
@@
Extension Study Across Different Modalities. Based on current experimental results, $\mathsf{PEFTGuard}$ performs well in detecting backdoor injections into LLMs using PEFT technology. Building on this, we further explore whether PEFTGuard is still effective in other modalities or Multimodal Large Language Models (MLLMs). Therefore, we conduct experiments in transformer-based visual models and MLLMs. Specifically, for the Visual Model (VM), we select Vit-base [13] as the base model and, following the experimental setup of BadNet [19], use a white $5\times 5$ pixel square in the bottom right corner of images from the CIFAR-10 [31] dataset as the trigger pattern. For the MLLM, we choose Qwen2-vl-2B [63] as the base model, inserting the trigger phrase “no cross, no crown” in the textual modality of the VQAv2 [18] dataset, with the target output set to “BOMB”. As shown in Table 7, $\mathsf{PEFTGuard}$ maintains high detection performance in VM as well, achieving a detection accuracy of $99.67\%$, and also reaches $99.33\%$ in MLLM. This demonstrates that $\mathsf{PEFTGuard}$ also works for other modalities or MLLMs.
@@
We also investigate the zero-shot transferability of $\mathsf{PEFTGuard}$ across several aspects, including PEFT methods, LoRA ranks, and attacks.
@@
Transferability on Different PEFT Methods. We aim to explore the zero-shot transferability of $\mathsf{PEFTGuard}$ on different PEFT methods. As indicated in Table 8, the classifier trained on the LoRA adapters is successfully transferred to QLoRA, LoRA+, and DoRA adapters, with detection accuracy exceeding $99\%$. Although $\mathsf{PEFTGuard}$ has demonstrated transfer capabilities across various PEFT methods, its transfer performance on AdaLoRA is weak. This discrepancy may be attributed to the fixed rank used in the LoRA, contrasted with AdaLoRA’s dynamic rank adjustment during training, which could result in different backdoor injection patterns and affect transferability.
@@
Transferability on Different Ranks of Adapters. We also investigate the zero-shot transferability of $\mathsf{PEFTGuard}$ across different ranks of adapters, which means using the classifier trained on LoRA adapters with rank $256$ transfer to other ranks under the same conditions. As detailed in Table 9, $\mathsf{PEFTGuard}$ exhibits excellent performance when transferred to datasets with LoRA ranks of less than $256$, achieving an average detection accuracy of more than $98.67\%$ and an average AUC of more than $0.998$. However, as the rank increases beyond $256$, the detection performance declines, with the lowest average accuracy recorded at $96.00\%$. This indicates that different LoRA ranks exhibit varying degrees of transferability.
Table 10: Evaluating Zero-Shot Transfer Performance Across Different Attacks. Blue indicates the training dataset.
@@
Transfer Dataset InsertSent $98.33\%$ / $0.996$ $50.00\%$ / $0.670$ $61.33\%$ / $0.845$ $50.00\%$ / $0.718$
RIPPLES $54.33\%$ /$0.851$ $100.00\%$ / $1.000$ $56.67\%$ / $0.828$ $50.00\%$ / $0.795$
Syntactic $54.33\%$ / $0.892$ $50.00\%$ / $0.722$ $99.33\%$ / $0.998$ $50.00\%$ / $0.849$
StyleBkd $64.33\%$ / $0.965$ $50.00\%$ / $0.662$ $95.33\%$ / $0.986$ $100.00\%$ / $1.000$
@@
Table 12: Performance metrics under various adaptive attacks methods. For the C&W attack, the detailed parameter settings are provided in Table 15.
Attack Method Parameters Backdoored Model Performance ASR on PEFTGuard

### v2

@@ Introduction
  • For the IMDB and AG News datasets, we apply four different textual backdoor attack methods. (1) RIPPLES [35] randomly inserts rare words from a dictionary as triggers to generate poisoned samples for backdoor training. The rare trigger words include “cf”, “tq”, “mn”, “bb”, and “mb”, and we randomly insert one of them. (2) InsertSent [11] uses a fixed sentence as a backdoor trigger, randomly inserting it into normal samples to generate poisoned samples. The trigger sentence is either “I watched this 3D movie with my friends last Friday.” (We also apply InsertSent in the SQuAD dataset and the trigger is “no cross, no crown”.) (3) Syntactic [54] modifies the sentence structures using SCPN [29], with the modified sentences serving as poisoned samples. The selected syntactic trigger template is $S(SBAR)(,)(NP)(VP)(.)$. (4) StyleBkd [53] uses a language model to convert the text’s style to another style, with the modified sentences used as poisoned samples. We choose the biblical style as the trigger.
@@
Other Hyperparameters. For all attacks, the poisoning rate is maintained at $5\%$. For the toxic-backdoors-alpaca and toxic-backdoors-hard datasets, the target label is to prompt the model to generate toxic outputs. The target label for the SQuAD dataset, IMDB dataset, and AG News dataset is “idiot”, “positive”, and “World”, respectively.
@@
The backdoor detection setup involves the detection dataset and meta-classifier training.
@@
Backdoor Detection Dataset. We leverage the same dataset generation process in Section 4.1 to generate a test dataset $\mathcal{D}_{\rm test}$. We use $\mathcal{D}_{\rm test}$ to evaluate the detection accuracy of $\mathsf{PEFTGuard}$. In terms of quantity, $|\mathcal{D}_{\rm train}|:|\mathcal{D}_{\rm test}|=8:2.$ We combine $\mathcal{D}_{\rm train}$ and $\mathcal{D}_{\rm test}$ into a single dataset, collectively named PADBench. Table 15 details the training and test sets for each task. Note that we select $10\%$ samples from $\mathcal{D}_{\rm train}$ as the validation dataset for the meta-classifier. Meanwhile, we highlight that in order to meet the practical application scenario, our PADBench includes adapters with different attack capabilities. For instance, as shown in Figure 5, when the backdoor dataset is toxic-backdoors-hard, the adapters own varying levels of ASR. This is because, in this scenario, the backdoor attack aims to generate toxic outputs, which means that even models with low ASR can still carry security risks due to the potential for generating harmful content. Therefore, we hope $\mathsf{PEFTGuard}$ can still catch the backdoored adapters even with low attack capability. Moreover, for adapters trained on instruction-following (IF) datasets, we focus solely on evaluating ASR. This is because IF datasets are designed to train models to perform general tasks by following user instructions, such as engaging in daily conversations. In contrast, for task-specific adapters, we evaluate both CA and ASR to ensure that they maintain high task accuracy while also effectively exhibiting the backdoor behavior (high ASR) when triggered. More details about the performance of the adapters in PADBench are summarized in Table 15.
Hyperparameters for Training Meta-Classifier. During the training phase of $\mathsf{PEFTGuard}_{\rm cls}$, we set the batch size to $4$, as well as using the Adam optimizer with an initial learning rate of $52e-$ 5 and a weight decay of $51e-$ 5. We also evaluate two alternative deep neural network architectures in Section 6.3, incorporating dropout layers in the networks with a dropout rate of $0.4$.
@@
Regarding backdoor mitigation, we consider three methods: SFT [55], DPO [56], and Fine-mixing [85] (More introductions are in Appendix B). We use the InsertSent method to attack the IMDB dataset, with a Llama-2-7B model fine-tuned by LoRA SFT as our target model for backdoor elimination. Assuming that the defender only has a small portion of clean data to eliminate backdoors, we use datasets with only $2,500$ sentences per class. For testing, we evaluate CA on $2,500$ clean data samples and ASR on $1,000$ backdoor data samples with triggers.
Evaluation
Based on the PADBench (Section 5.3) and the experimental settings (Section 5.3.2), we conduct a systematic evaluation of $\mathsf{PEFTGuard}$, which includes its detection performance on various datasets and attacks, comparison with other SOTA backdoor detection methods, efficacy across different PEFT methods, detection capabilities on various base models, performance across different target projection matrices of adapters, and effectiveness under different training quantity. In addition, we assess the transferability of $\mathsf{PEFTGuard}$ and its robustness against adaptive attacks. Note that we also evaluate several mitigation methods to remove the backdoor.
@@
As shown in Table 1, $\mathsf{PEFTGuard}$ effectively detects backdoors for IF datasets, specifically achieving detection accuracy of $100.00\%$ and detection AUC of $1.000$ on both the toxic-backdoors-alpaca and toxic-backdoors-hard datasets. When confronted with various attacks on the topic classification dataset (AG News), $\mathsf{PEFTGuard}$ achieves $100.00\%$ accuracy and $1.000$ AUC under RIPPLES and StyleBkd attacks. For InsertSent/Syntactic attacks, it achieves $98.33\%$/$99.33\%$ accuracy and $0.996$/$0.998$ AUC. Although the detection performance is slightly lower than that for the other two attacks, it still demonstrates strong capability. Similarly, various backdoor attacks applied to the sentiment classification dataset IMDB also demonstrate the robust performance of $\mathsf{PEFTGuard}$. These results indicate that $\mathsf{PEFTGuard}$ consistently maintains high detection performance across different datasets and textual backdoor attack methods, demonstrating its stability and effectiveness.
@@
Baselines. We consider four SOTA detection baselines: (1) Trojan-Miner [2] trains a seq-to-seq model to detect classifiers potentially containing backdoors and to generate text sequences that may include parts or all of a Trojan trigger. (2) AttenTD [44] uses a set of neutral trigger candidates and attention anomalies to distinguish models infected with backdoors. (3) PICCOLE [41] uses optimization to invert the distribution of words to indicate their likelihood in triggers, utilizing discriminative analysis of words to determine if the model is particularly sensitive to potential trigger words. (4) MNTD [70] employs a query set and meta-training with the representation (hidden state of the last layer) obtained from the detection model while optimizing both the query set and training the meta-classifier. Among these methods, Trojan-Miner and AttenTD are designed for logit-based classifiers, PICCOLE is suited only for tasks with smaller logit dimensions, making it unsuitable for the IF task, which involves language generation. Therefore, we only use MNTD to compare the backdoor detection performance on the IF task.
Attack Scenarios. Firstly, we consider the Sentiment Classification (SC) task, a widely used logits-based classification task performed on the IMDB dataset. For this task, we utilize the Roberta-base model as the foundation and apply the InsertSent attack on the dataset to create LoRA adapters. Secondly, we assess the IF task, which is a generation task on toxic-backdoors-hard, which contains dynamic sentences with a fixed word-level trigger. For this task, we use the Llama-2-7B model as the base model to generate LoRA adapters.
Results Analysis. For each detection method, we compare the best results achieved. As shown in Table 2, $\mathsf{PEFTGuard}$ reaches a detection accuracy of $98.33\%$ and an AUC of $1.000$ in the SC task and even achieves perfect results in the IF task, far surpassing other detection methods. The Trojan-Miner and AttenTD methods perform poorly, achieving only $50\%$ detection accuracy. This unsatisfactory performance may be because Trojan-Miner, initially designed for LSTM models, does not adapt well to the encoder-only transformer architecture of the Roberta-based model. Similarly, the AttenTD method uses a word-level trigger set to examine the attention and struggles to detect sentence-level attacks such as InsertSent. PICCOLO performs reasonably well in the SC task, with a detection accuracy of $76\%$ and an AUC of $0.890$, indicating the moderate ability to differentiate models. Although MNTD performs well in the SC task, achieving $88\%$ accuracy and $0.937$ AUC, it performs poorly in the IF task, with only $51\%$ accuracy and $0.510$ AUC. In general, we consider $\mathsf{PEFTGuard}$ as the best backdoor detection method since it outperforms other detection methods in both tasks.
@@
Different PEFT Methods. Besides LoRA, there are different PEFT methods. In this part, we aim to evaluate the performance of $\mathsf{PEFTGuard}$ across these PEFT-based adapters. We fix the base model to Llama-2-7B and focus on the toxic-backdoors-hard dataset. As shown in Table 3, $\mathsf{PEFTGuard}$ achieves excellent detection results for these five different PEFT methods. Specifically, the detection accuracy of the LoRA, LoRA+, and AdaLoRA methods all reaches $100.00\%$ with an AUC of $1.000$. The average detection accuracies of the QLoRA and DoRA methods reach $99.67\%$ and $98.00\%$, respectively, with both methods achieving an AUC of $1.000$. These results indicate that $\mathsf{PEFTGuard}$ is effective in backdoor detection on models trained with different PEFT methods.
Table 4: Effectiveness of $\mathsf{PEFTGuard}$ on different transformer-based architecture models.
@@
Different Base Models. We then investigate if $\mathsf{PEFTGuard}$ is effective for different base models. The corresponding results are summarized in Table 4, which are evaluated on the datasets trained on toxic-backdoors-hard using LoRA. We observe that, for all base models, the backdoor detection AUC is $1.000$. Specifically, for Llama-3-8B, Llama-2-7B, Qwen1.5-7B-Chat, and Flan-t5-xl adapters, the detection accuracy reaches $100\%$. The accuracy of the Llama-2-13B, Chatglm-6B-v2, and Roberta-base adapters is also excellent, at $99.67\%$, $99.33\%$, and $98.33\%$, respectively. In summary, we have the following insights.
@@
    Based on the performances across the Llama series (with model sizes of 7B, 8B, and 13B), we demonstrate the effectiveness of $\mathsf{PEFTGuard}$ on different parameter scales.
@@
    As mentioned in Section 5.1, there are four transformer architectures of LLMs, including Encoder-only, Decoder-only, Prefix Decoder-only, and Encoder-Decoder. We also consider the three most common transformer attention mechanisms: Multi-Head Attention, Grouped Query Attention, and Multi-Query Attention. $\mathsf{PEFTGuard}$ demonstrates general performances across these model architectures and attention mechanisms, indicating its effectiveness for backdoor detection in transformer-based LLMs.
Table 5: Effectiveness of $\mathsf{PEFTGuard}$ on different target projection matrices.
@@
Different Target Projection Matrices. As discussed in Section 4.2, we fine-tune adapters primarily by focusing on $\Delta_{q}$ and $\Delta_{v}$. However, it is equally important to explore whether $\mathsf{PEFTGuard}$ can be applied to other weight matrices within the self-attention module, as these matrices have also proven to be effective in tuning strategies [25]. For instance, in adapter training on $\Delta_{q}$, $\Delta_{k}$, $\Delta_{v}$, and $\Delta_{o}$, the concatenation of matrices in the $l$-th layer can be represented as $\Delta^{(l)}_{\rm concat}=[\Delta^{(l)}_{q},\Delta^{(l)}_{k},\Delta^{(l)}_{v},\Delta^{(l)}_{o}]$.
Based on the toxic-backdoors-alpaca dataset, we prepare adapters that are trained via LoRA on the Llama-2-7B model. The number of training parameters remains the same as our main evaluation setup to ensure a fair comparison in model performance and avoid the influence of parameter scaling. As shown in Table 5, our $\mathsf{PEFTGuard}$ demonstrates robust performance across various target projection matrices of adapters, all achieving $100\%$ detection accuracy.
@@
Impact of Training Quantities. As mentioned in Section 5.3.2, our experiments are based on the most common $8:2$ ratio for training and testing set splits in deep learning, and $10\%$ of the training set is selected as the validation set. Here we also investigate the performance of $\mathsf{PEFTGuard}$ under different ratios of training datasets. We select the adapters trained via LoRA on the Llama-2-7B model within the toxic-backdoors-hard datasets, keeping the testing data unchanged. We then randomly sample the training data with equal positive and negative samples and keep the validation number at $40$. As shown in Figure 6, $\mathsf{PEFTGuard}$ performs well across different data quantities ranging from $100$ to $360$. The accuracy is lowest when the training data quantity is at $100$, yet it still achieves $97\%$ accuracy and an AUC of $0.995$. When the training data quantity increases to $200$, $\mathsf{PEFTGuard}$ consistently reaches $100\%$ accuracy.
Table 6: Performance of $\mathsf{PEFTGuard}$ with different architectures. (AdaptiveAvgPool refers to replacing the Convolutional Layer in Figure 6(a) with an Adaptive Average Pooling Layer, while MaxPool indicates replacement with a Max Pooling Layer.)
@@
Explorations of $\mathsf{PEFTGuard}$ Architecture. Referring to Section 4.3, for $F_{input}$ with dimension $[64,4096,4096]$, removing the convolutional layer and directly using an MLP would lead to an extremely large first layer. This occurs because the high-dimensional input lacks prior reduction, vastly increasing parameters and making the classifier hard to train and implement. Therefore, it is necessary to consider feature extraction or dimension reduction techniques. Here we first discuss the ablation study of dimensionality reduction strategy, where we replace the convolutional layer with a pooling layer. Then we explore other efficient model architectures, demonstrating that the $\mathsf{PEFTGuard}$ framework is broadly adaptable and not limited to the default structure shown in Figure 6(a).
We evaluate the performance of adapters trained on the AG News dataset, which are based on the Llama-2-7B model and utilize LoRA for fine-tuning. After replacing the convolutional layer with adaptive pooling layers, the detection performance is weaker, achieving only a $50\%$ accuracy. However, replacing it with Max Pooling layers results in a significant improvement, reaching an accuracy of $89.33\%$.
To explore alternative structures for $\mathsf{PEFTGuard}$ classifier, we first increase the number of convolutional layers, incorporate adaptive pooling layers, and reduce the number of fully connected layers, resulting in architecture A1. Building on A1, we then replace the intermediate convolutional layers with residual layers, forming architecture A2 (refer to Figure 6(b)). The comparable performance indicates that the original architecture is a suitable choice.
Table 7: Effectiveness of $\mathsf{PEFTGuard}$ on various modality models.
@@
Extension Study Across Different Modalities. Based on current experimental results, $\mathsf{PEFTGuard}$ performs well in detecting backdoor injections into LLMs using PEFT technology. Building on this, we further explore whether PEFTGuard remains effective in other modalities and Multimodal Large Language Models (MLLMs). Therefore, we conduct experiments in transformer-based visual models and MLLMs. Specifically, for the Visual Model (VM), we select Vit-base [15] as the base model and, following the experimental setup of BadNets [21], use a white $5\times 5$ pixel square in the bottom right corner of images from the CIFAR-10 [33] dataset as the trigger pattern. For the MLLM, we choose Qwen2-vl-2B [68] as the base model, inserting the trigger phrase “no cross, no crown” in the textual modality of the VQAv2 [20] dataset, with the target output set to “BOMB”. As shown in Table 7, $\mathsf{PEFTGuard}$ maintains high detection performance in VM as well, achieving a detection accuracy of $99.67\%$, with $99.33\%$ in MLLM. This demonstrates that $\mathsf{PEFTGuard}$ also works for other modalities or MLLMs.
@@
We further investigate the zero-shot transferability of $\mathsf{PEFTGuard}$ across multiple aspects, including PEFT methods, LoRA ranks, and attacks.
@@
Transferability on Different PEFT Methods. We aim to explore the zero-shot transferability of $\mathsf{PEFTGuard}$ on different PEFT methods. As indicated in Table 8, the classifier trained on the LoRA is successfully transferred to QLoRA, LoRA+, and DoRA adapters, with detection accuracy exceeding $99\%$. While $\mathsf{PEFTGuard}$ demonstrates transferability across various PEFT methods, its transfer performance on AdaLoRA is weak. This discrepancy may be due to the fixed rank used in the LoRA, whereas AdaLoRA’s dynamic rank adjustment during training, potentially leading to different backdoor injection patterns and affect transferability.
@@
Transferability on Different Ranks of Adapters. We also investigate the zero-shot transferability of $\mathsf{PEFTGuard}$ across different ranks of adapters, meaning that using the classifier trained on LoRA adapters with rank $256$ is transferred to other ranks under the same conditions. Table 9 shows that $\mathsf{PEFTGuard}$ exhibits excellent performance when transferred to LoRA ranks of less than $256$, achieving an average detection accuracy of more than $98.67\%$ and an average AUC of more than $0.998$. However, as the rank exceeds $256$, the detection performance declines, with the lowest average accuracy recorded at $96.00\%$. This highlights the impact of LoRA rank on transferability.
Table 10: Zero-Shot transferability across different attacks. Blue indicates the training dataset.
@@
Transfer Dataset InsertSent $\mathbf{98.33\%}$ /$\mathbf{0.996}$ $50.00\%$ /$0.670$ $61.33\%$ /$0.845$ $50.00\%$ /$0.718$
RIPPLES $54.33\%$ /$0.851$ $\mathbf{100.00\%}$ /$\mathbf{1.000}$ $56.67\%$ /$0.828$ $50.00\%$ /$0.795$
Syntactic $54.33\%$ /$0.892$ $50.00\%$ /$0.722$ $\mathbf{99.33\%}$ /$\mathbf{0.998}$ $50.00\%$ /$0.849$
StyleBkd $64.33\%$ /$0.965$ $50.00\%$ /$0.662$ $95.33\%$ /$0.986$ $\mathbf{100.00\%}$ /$\mathbf{1.000}$
@@
Table 12: Performance metrics under various adaptive attacks methods. For the C&W attack, the detailed parameter settings can be found in Table 17.
Attack Method Parameters Performance of Backdoored Model ASR on PEFTGuard